# An Explainable Framework for Particle Swarm Optimization using Landscape Analysis and Machine Learning

### v1

@@ 3 Explainablity of PSO
3.1 Population Topologies
@@ 4 Experimental Setup and Results
Our experimental evaluation examines all twenty four BBOB functions across two problem dimensions (2d and 5d) for each PSO configuration. The computational budget is set at 100 and 500 iterations, with three distinct swarm topologies Star, Ring, and Von Neumann. The baseline PSO parameters for each topology in both dimensional spaces are detailed in Table 7. Our methodology for ensuring statistical reliability involves conducting five independent runs across the first five instances of each benchmark function. Algorithm performance is quantified using the normalized Area Over the Convergence Curve (AOCC) (Eq. (3)), which provides a comprehensive measure of optimization efficiency by integrating solution quality across the entire search process [50]. This metric enables direct comparison of different configurations’ effectiveness throughout the optimization trajectory.
@@ 5 Hyperparameter Impact Analysis
We examined $1,728$ PSO configurations across twenty four BBOB functions, focusing on the impact of different communication topologies and hyperparameters on performance. Using SHAP-based plots for 2d & 5d (Table 7), we analyse the contribution parameters such as n_particles, $c_{1}$, $c_{2}$, $p$, and $k$. The SHAP values indicate how each parameter affects the AOCC, with yellow dots representing positive contributions and violet dots indicating negative ones. There is a clear distinction between the two colours across all topologies and parameters, such as n_particles, $c_{1}$, $c_{2}$, $p$, $k$, etc. and others. Specifically, yellow dots represent positive SHAP values, while violet dots consistently indicate negative SHAP values.
@@ 5.3 PSO Variant configuration selection
The “single-best mean” (sbm) and “single-best std” (sbs) are the statistical measures used to evaluate the consistency and performance of the global best solution ($g_{best}$) over multiple independent runs. The “sbm” represents the average of the best solutions found across all runs, indicating the algorithm’s overall effectiveness in locating near-optimal solutions. The “sbs” measures the variability in these solutions a low standard deviation suggests reliable convergence, while a high deviation indicates instability or sensitivity to initial conditions. Similarly, “avg-best mean” (abm) and “avg-best std” (abs) assess the collective behaviour of the swarm by analysing the average of particles’ personal best positions ($p_{best,i}$) over multiple runs. The abm reflects the typical performance level of the swarm, while the abs quantifies how consistently the swarm performs across different runs. A low abs implies stable exploration, whereas a high value suggests erratic convergence or poor swarm coordination. Together, these metrics help fine-tune PSO parameters, ensuring robust optimization by balancing exploration and exploitation.
@@
Similarly, the results in Table 9 For unimodal functions, such as $f_{12}$, all topologies perform exceptionally well, achieving near-optimal results (e.g., means of $2.00\times 10^{-3}$ ) with negligible variability (standard deviations as low as $8.90\times 10^{-19}$), indicating robustness in simple landscapes. However, in multi-modal functions like $f_{4}$ and $f_{6}$, differences emerge. The Ring topology often excels in maintaining diversity, as seen in $f_{4}$ with a lower single-best standard deviation ($6.21\times 10^{-3}$) compared to the Star topology ($1.00\times 10^{-2}$). Meanwhile, the Von Neumann topology strikes a balance between exploration and exploitation, performing well in functions like $f_{6}$ with a high $R^{2}$ ratio ($9.70\times 10^{-1}$), significantly outperforming the Star topology ($8.84\times 10^{-3}$). For highly complex or deceptive functions, such as $f_{7}$ and $f_{14}$, the Von Neumann topology demonstrates greater resilience. For instance, in $f_{14}$, it achieves an overall mean of $1.81\times 10^{-2}$, while the Ring topology occasionally dominates in specific cases like $f_{7}$ with a $R^{2}$ ratio of $9.12\times 10^{-1}$. The Star topology, though fast in convergence, exhibits higher variability and inconsistency, as seen in $f_{7}$ with a single-best standard deviation of $2.75\times 10^{-2}$ and in $f_{9}$ with erratic performance despite a low overall mean ($1.00\times 10^{-2}$ ).
@@ 5.3.1 Time Complexity Analysis
The impact of swarm topology on PSO’s computational efficiency becomes increasingly significant when examining different problem dimensions. For 2-dimensional problems, our experiments reveal clear differences in execution times: the Von Neumann topology demonstrates superior efficiency (6.56 hours), followed by Star (7.2 hours) and Ring (9.35 hours). This pattern persists and amplifies in 5-dimensional problems, with Von Neumann maintaining its advantage (56.87 hours), while Star and Ring require 61.17 hours and 75.52 hours, respectively. The consistent performance gap across dimensions suggests that Von Neumann’s structured yet decentralized communication pattern scales more effectively, likely due to its balanced information flow that avoids bottlenecks. In contrast, the Ring topology’s sequential neighbor-based updates appear to compound its time complexity as dimensionality increases, resulting in significantly longer runtimes. The Star topology occupies an intermediate position, with its global communication structure offering better scalability than Ring but still incurring higher costs than Von Neumann. These findings underscore how topology selection becomes increasingly critical for computational efficiency in higher-dimensional optimization problems, where the differences in time complexity grow substantially. The results provide practical guidance for algorithm design, particularly in scenarios where both solution quality and computational resources must be carefully balanced.
@@ 6 Data-Driven and Interpretable Configuration Learning
In the IOHxplainer pipeline, extensive data is gathered on the behaviour and performance of various algorithm configurations for each topologies. While explainable AI (XAI) techniques can extract key insights from this data, there are further advantages to systematically storing and re-using it. In this section, we highlight several practical ways in which the collected data can be used to enhance the selection, configuration, and understanding of algorithm configurations.
@@
The IOHxplainer toolbox offers a valuable secondary benefit: it enables the straightforward training of machine learning models that can predict the most suitable algorithm configuration including both the choice of modules and their hyper-parameters based on observed characteristics of the optimization landscape. Figure 8 represents the workflow for performance classification using Exploratory Landscape Analysis (ELA) features. The diagram illustrates the process from feature extraction and augmentation to model training and evaluation for predicting PSO algorithm performance based on AOCC metrics. This is essentially the task of automated algorithm configuration and selection, echoing the goals of previous research such as those referenced in [24, 30] and [48]. What distinguishes our approach is that we merge the selection of the algorithm and the tuning of its configuration into a single, unified machine learning problem. We achieve this by employing multi-output models that can handle both classification (for discrete choices like module selection) and regression (for continuous settings like hyper-parameters) simultaneously. Specifically, our models are restricted to shallow Decision Tree (DT) [46] and Random Forest (RF) [39], chosen for their simplicity and interpretability. We gathered a wide range of Exploratory Landscape Analysis (ELA) features to show what we can do. To create these features, we set up an experiment where we sampled 1000 points from each instance in the BBOB suite. For each instance, we aimed for our machine learning models to find the best-performing algorithm configuration from our experimental results. Using models like RF and easy-to-understand DT, we can train our systems to predict which algorithm configurations will perform well based on the ELA features. This approach not only automates the process of algorithm selection and configuration, but also provides transparent and understandable decision rules thanks to the interpretable nature of the models used.
@@ 7 Conclusion
This work makes three key contributions to explainablity of PSO algorithm. First, we demonstrate the critical importance of landscape analysis through Exploratory Landscape Analysis (ELA), which provides fundamental insights into problem structure, multimodality, and difficulty characteristics. Our work shows how ELA features can guide algorithm selection and configuration by revealing intrinsic properties of optimization landscapes that influence heuristic performance. The IOHxplainer framework significantly enhances this analysis by enabling systematic investigation of hyperparameter contributions, particularly in quantifying how different PSO topologies interact with specific landscape features. Second, we apply this explainable approach to conduct the most comprehensive analysis of PSO topologies such as Star, Ring and Von Neumann configurations. The redesigned IOHxplainer toolbox turns out to be an invaluable asset in identifying how topological structure affects swarm behaviour, with Von Neumann consistently demonstrating superior time efficiency while Ring topology better maintains diversity in complex landscapes. Finally, we introduce a novel machine learning application by training Random Forest and Decision Tree classifiers on AOCC data. These models successfully predict optimal topology selection based on landscape characteristics, achieving accuracy. While focused on PSO, our methodology establishes a template for explainable analysis of any iterative swarm based meta-heuristic. Future directions include extending the framework to higher dimensions, integrating additional XAI techniques, and developing adaptive topology selection systems. This work advances optimization research from empirical benchmarking toward truly explainable algorithm design, with implications for both theoretical understanding and real-world applications.
@@
Nitin Gupta:
Writing – original draft, Methodology, Experimentation. Bapi Dutta: Review & editing. Anupam Yadav: Writing – review & editing, Supervision.

### v2

@@ 3 Explainablity of PSO
3.1 Swarm Behaviour Analysis on Population Topologies
@@ 4 Experimental Setup and Results
Our experimental evaluation examines all twenty four BBOB functions across two problem dimensions (2d and 5d) for each PSO configuration. The computational budget is set at 100 and 500 iterations, with three distinct swarm topologies Star, Ring, and Von Neumann. The baseline PSO parameters for each topology in both dimensional spaces are detailed in Table 7. Our methodology for ensuring statistical reliability involves conducting five independent runs across the first five instances of each benchmark function.
4.1 Performance Metrics
To evaluate the generated algorithms effectively over a complete set of benchmark functions, we use an anytime performance measure. This approach is critical because it captures the entire convergence profile of an algorithm, thereby rewarding those that find good solutions quickly and reliably, not just those that converge to a decent solution by the final budget. Specifically, it quantifies performance over the complete budget, instead of only looking at the final objective function value. We use the normalized Area Over the Convergence Curve (AOCC) (Eq. (3)), as introduced in  [55]. The AOCC is calculated as:
@@ 5 Hyperparameter Impact Analysis
We examined $1,728$ PSO configurations across twenty four BBOB functions, focusing on the impact of different communication topologies and hyperparameters on performance. Using SHAP-based plots for 2d & 5d (Table 7), we analyse the contribution parameters such as n_particles, $c_{1}$, $c_{2}$, $w$, $p$, and $k$. The hyperparameters for the PSO algorithm were selected based on established conventions and empirical ranges from the literature. The cognitive ($c_{1}$) and social ($c_{2}$) coefficients were chosen from intervals that adhere to the common stability criterion of ($c_{1}+c_{2}<4$), which helps prevent particle velocity divergence and promotes effective convergence, as analyzed by [10]. The inertia weight ($w$) values of {0.9, 0.5, 0.7} cover standard strategies, where a high value (0.9) encourages global exploration by maintaining momentum, and lower values (0.5, 0.7) facilitate local exploitation, following the principles introduced by [48]. Swarm sizes ($n$) of {50, 100, 150} were selected as a representative range commonly used in the literature, as surveyed by [6]. Finally, the parameters for neighborhood topology ($k,p,r$) were limited to low integer values to model fundamental local interaction structures—such as the Ring and Von Neumann topologies identified as high-performing by [24]—without inducing excessive computational complexity. The SHAP values indicate how each parameter affects the AOCC, with yellow dots representing positive contributions and violet dots indicating negative ones. There is a clear distinction between the two colours across all topologies and parameters, such as n_particles, $c_{1}$, $c_{2}$, $p$, $k$, etc. and others. Specifically, yellow dots represent positive SHAP values, while violet dots consistently indicate negative SHAP values.
@@ 5.3 PSO Variant configuration selection
This analysis distinguishes between specialized and general hyperparameter configurations using four key metrics. The “Single Best Mean” (sbm) and “Single Best Std” (sbs) quantify the performance of a configuration optimized for a single function; the mean represents its average performance score, while the standard deviation captures the variability of that performance across different algorithm instances and random seeds. Conversely, the “Avg-Best Mean” (abm) and “Avg-Best Std” (abs) describe the best general-purpose configuration, which is selected for its robust performance across all benchmark functions. Here, the abm indicates its overall average performance, and the abs measures the consistency of that performance when applied to the diverse set of functions. In essence, the “Single Best” metrics reflect specialized, peak performance on a specific function, whereas the “Avg-Best” metrics reflect the reliability and generalizability of a configuration across a broader problem domain.
@@
Similarly, the results in Table 9 For unimodal functions, such as $f_{12}$, all topologies perform exceptionally well, achieving near-optimal results (e.g., means of $2.00\times 10^{-3}$ ) with negligible variability (standard deviations as low as $8.90\times 10^{-19}$), indicating robustness in simple landscapes. However, in multi-modal functions like $f_{4}$ and $f_{6}$, differences emerge. The Ring topology often excels in maintaining diversity, as seen in $f_{4}$ with a lower single-best standard deviation ($6.21\times 10^{-3}$) compared to the Star topology ($1.00\times 10^{-2}$). Meanwhile, the Von Neumann topology strikes a balance between exploration and exploitation, performing well in functions like $f_{6}$ with a high $R^{2}$ ratio ($9.70\times 10^{-1}$), significantly outperforming the Star topology ($8.84\times 10^{-3}$). For highly complex or deceptive functions, such as $f_{7}$ and $f_{14}$, the Von Neumann topology demonstrates greater resilience. For instance, in $f_{14}$, it achieves an overall mean of $1.81\times 10^{-2}$, while the Ring topology occasionally dominates in specific cases like $f_{7}$ with a $R^{2}$ ratio of $9.12\times 10^{-1}$. The Star topology, though fast in convergence, exhibits higher variability and inconsistency, as seen in $f_{7}$ with a single-best standard deviation of $2.75\times 10^{-2}$ and in $f_{9}$ with erratic performance despite a low overall mean ($1.00\times 10^{-2}$ ). The optimal parameter configurations summarized in Table 10 provide evidence for the specialized roles of different population topologies. The prevalence of a high moment of inertia ($w$ = 0.9) in the Ring topology indicates its primary strength lies in exploration. The structural constraint of the ring, with its low connectivity, naturally slows information flow; a high inertia value compensates for this by sustaining particle velocity, thus preventing premature convergence. On the other hand, the frequent selection of a moderate inertia (w = 0.5) for the highly connected Star and Von Neumann topologies aligns with their aptitude for exploitation. The rapid information dissemination in these structures quickly identifies promising regions, and a lower inertia allows the swarm to fine-tune its search and converge efficiently within those regions.
@@ 5.3.1 Time Complexity Analysis
The impact of swarm topology on PSO’s computational efficiency becomes increasingly significant when examining different problem dimensions. For 2-dimensional problems, our experiments reveal clear differences in execution times: the Von Neumann topology demonstrates superior efficiency (6.56 hours), followed by Star (7.2 hours) and Ring (9.35 hours). This pattern persists and amplifies in 5-dimensional problems, with Von Neumann maintaining its advantage (56.87 hours), while Star and Ring require 61.17 hours and 75.52 hours, respectively. The consistent performance gap across dimensions suggests that Von Neumann’s structured yet decentralized communication pattern scales more effectively, likely due to its balanced information flow that avoids bottlenecks. In contrast, the Ring topology’s sequential neighbor-based updates appear to compound its time complexity as dimensionality increases, resulting in significantly longer runtimes. The Star topology occupies an intermediate position, with its global communication structure offering better scalability than Ring but still incurring higher costs than Von Neumann. These findings underscore how topology selection becomes increasingly critical for computational efficiency in higher-dimensional optimization problems, where the differences in time complexity grow substantially. The results provide practical guidance for algorithm design, particularly in scenarios where both solution quality and computational resources must be carefully balanced. The computational demand of this experiment is profoundly high; a single function evaluation on a 20-dimensional problem already requires approximately 5 hours to complete 30 repetitions, making the process extremely time-consuming. The experimental results on 20-dimensional problems have been uploaded to the provided GitHub link. This heavy burden highlights the necessity for high-core CPU architectures to parallelize evaluations and manage the immense processing load. Unfortunately, the prohibitive computational power required to extend this study to even higher dimensions or a broader set of test functions is currently beyond our available resources, thereby limiting the scope of our experimental analysis.
@@
Table 10: Single best Config of Star, Ring and Von Neumann Topologies over all BBOB functions on d=5.
Function Topologies $c_{1}$ $c_{2}$ $k$ n_particlen\_particle $p$ $r$ $w$
f1 Star 0.9 0.7 1 50 1 1 0.5
Ring 0.3 0.7 3 50 2 1 0.9
Von Neumann 0.9 0.7 1 50 1 1 0.5
f2 Star 0.7 0.6 1 50 1 1 0.5
Ring 0.5 0.6 3 50 2 1 0.7
Von Neumann 0.3 0.6 1 50 1 1 0.5
f3 Star 0.9 0.7 1 50 1 1 0.5
Ring 0.5 0.7 3 50 1 1 0.7
Von Neumann 0.9 0.7 1 50 1 1 0.5
f4 Star 0.7 0.7 1 50 1 1 0.5
Ring 0.5 0.7 3 50 2 1 0.9
Von Neumann 0.3 0.6 1 50 1 1 0.5
f5 Star 0.7 0.7 1 50 1 1 0.5
Ring 0.5 0.2 1 100 1 1 0.9
Von Neumann 0.5 0.7 1 50 1 1 0.5
f6 Star 0.5 0.7 1 50 1 1 0.5
Ring 0.7 0.7 3 50 2 1 0.9
Von Neumann 0.9 0.6 1 50 1 1 0.5
f7 Star 0.9 0.6 1 50 1 1 0.5
Ring 0.3 0.6 3 50 1 1 0.9
Von Neumann 0.7 0.7 1 50 1 1 0.5
f8 Star 0.9 0.6 1 50 1 1 0.5
Ring 0.9 0.6 3 50 1 1 0.7
Von Neumann 0.5 0.7 1 50 1 1 0.5
f9 Star 0.9 0.6 1 50 1 1 0.5
Ring 0.9 0.7 3 50 2 1 0.5
Von Neumann 0.9 0.6 1 50 1 1 0.5
f10 Star 0.7 0.4 1 50 1 1 0.7
Ring 0.9 0.6 3 50 1 1 0.9
Von Neumann 0.5 0.6 1 50 1 1 0.5
f11 Star 0.9 0.4 1 50 1 1 0.7
Ring 0.9 0.7 3 50 2 1 0.9
Von Neumann 0.7 0.6 1 50 1 1 0.7
f12 Star 0.7 0.6 1 50 1 1 0.5
Ring 0.3 0.6 1 50 1 1 0.7
Von Neumann 0.7 0.6 1 50 1 1 0.5
f13 Star 0.7 0.6 1 50 1 1 0.5
Ring 0.3 0.7 3 50 2 1 0.9
Von Neumann 0.7 0.7 1 50 1 1 0.5
f14 Star 0.7 0.7 1 50 1 1 0.5
Ring 0.5 0.6 3 150 1 1 0.9
Von Neumann 0.5 0.7 1 50 1 1 0.5
f15 Star 0.5 0.6 1 50 1 1 0.5
Ring 0.7 0.6 3 50 1 1 0.9
Von Neumann 0.9 0.6 1 50 1 1 0.5
f16 Star 0.5 0.7 1 50 1 1 0.5
Ring 0.9 0.4 3 50 2 1 0.7
Von Neumann 0.9 0.7 1 50 1 1 0.5
f17 Star 0.7 0.6 1 50 1 1 0.5
Ring 0.3 0.7 3 50 2 1 0.7
Von Neumann 0.9 0.6 1 50 1 1 0.5
f18 Star 0.7 0.6 1 50 1 1 0.5
Ring 0.3 0.7 3 50 2 1 0.7
Von Neumann 0.5 0.7 1 50 1 1 0.5
f19 Star 0.3 0.4 1 50 1 1 0.5
Ring 0.5 0.6 3 50 1 1 0.9
Von Neumann 0.9 0.6 1 50 1 1 0.7
f20 Star 0.9 0.7 1 50 1 1 0.5
Ring 0.3 0.7 3 50 2 1 0.9
Von Neumann 0.9 0.7 1 50 1 1 0.5
f21 Star 0.3 0.7 1 100 1 1 0.5
Ring 0.5 0.4 3 50 1 1 0.7
Von Neumann 0.3 0.7 1 50 1 1 0.5
f22 Star 0.3 0.6 1 50 1 1 0.5
Ring 0.5 0.7 3 50 1 1 0.7
Von Neumann 0.9 0.7 1 50 1 1 0.5
f23 Star 0.9 0.7 1 50 1 1 0.7
Ring 0.7 0.6 3 100 2 1 0.9
Von Neumann 0.9 0.6 1 150 1 1 0.5
f24 Star 0.9 0.7 1 50 1 1 0.5
Ring 0.7 0.6 3 150 2 1 0.9
Von Neumann 0.5 0.7 1 150 1 1 0.7
@@ 6 Data-Driven and Interpretable Configuration Learning
In the IOHxplainer pipeline, extensive data is gathered on the behaviour and performance of various algorithm configurations for each topologies. While explainable AI (XAI) [11] techniques can extract key insights from this data, there are further advantages to systematically storing and re-using it. In this section, we highlight several practical ways in which the collected data can be used to enhance the selection, configuration, and understanding of algorithm configurations.
@@
The IOHxplainer toolbox offers a valuable secondary benefit: it enables the straightforward training of machine learning models that can predict the most suitable algorithm configuration including both the choice of modules and their hyper-parameters based on observed characteristics of the optimization landscape. Figure 8 represents the workflow for performance classification using Exploratory Landscape Analysis (ELA) features. The diagram illustrates the process from feature extraction and augmentation to model training and evaluation for predicting PSO algorithm performance based on AOCC metrics. This is essentially the task of automated algorithm configuration and selection, echoing the goals of previous research such as those referenced in [26, 33] and [53]. What distinguishes our approach is that we merge the selection of the algorithm and the tuning of its configuration into a single, unified machine learning problem. We achieve this by employing multi-output models that can handle both classification (for discrete choices like module selection) and regression (for continuous settings like hyper-parameters) simultaneously. Specifically, our models are restricted to shallow Decision Tree (DT) [51, 57] and Random Forest (RF) [44]. This choice was strategic, balancing the competing needs of interpretability and predictive performance. The shallow DT was chosen for its role as a transparent, ‘white-box’ model; its limited depth yields a simple set of human-readable decision rules that directly link Exploratory Landscape Analysis (ELA) features to algorithm performance, providing crucial insights for domain experts. To complement this and ensure high accuracy, we employed a RF model. As an ensemble method, RF mitigates the potential overfitting of individual trees and robustly captures the complex, non-linear relationships between the wide range of ELA features and the optimal algorithm configuration. We gathered a wide range of Exploratory Landscape Analysis (ELA) features to show what we can do. To create these features, we set up an experiment where we sampled 1000 points from each instance in the BBOB suite. For each instance, we aimed for our machine learning models to find the best-performing algorithm configuration from our experimental results. Using this two-pronged modeling approach-RF for powerful predictions and DT for transparent rule extraction-we can train our systems to predict which algorithm configurations will perform well based on the ELA features. This approach not only automates the process of algorithm selection and configuration, but also provides transparent and understandable decision rules thanks to the interpretable nature of the models used.
@@ 6.1 Explaining PSO via Decision Trees and ELA Features
6.2 Limitations
A primary limitation of this experimental setup is its expensive computational intensity and scalability. The methodology is not configurable for GPU acceleration and instead demands substantial CPU resources, requiring a high number of cores for parallel evaluation. This results in a considerable time complexity, as the exhaustive sampling of hyperparameter configurations and multiple algorithm runs on various benchmark functions is inherently time-consuming. Consequently, our current computational power imposed a strict ceiling on the scale of the experiments, preventing an extension to higher-dimensional problems, a broader set of functions, or a more fine-grained hyperparameter search, which would be necessary for even more comprehensive insights.
@@ 7 Conclusion
This work advances the explainability of the Particle Swarm Optimization (PSO) algorithm through three key contributions. First, it highlights the importance of Exploratory Landscape Analysis (ELA) in revealing problem structure, multimodality, and landscape difficulty, demonstrating how these features can guide algorithm selection and configuration. The IOHxplainer framework further enables systematic assessment of hyperparameter effects, particularly elucidating interactions between PSO topologies and landscape characteristics. Second, the work presents a comprehensive evaluation of Star, Ring, and Von Neumann topologies. Results indicate that Star topology excels on unimodal or smooth functions with rapid convergence and high stability; Ring topology performs best on multimodal landscapes, balancing exploration and exploitation; and Von Neumann topology offers robust, general-purpose performance. These findings provide actionable insights into topology selection based on landscape properties. Finally, a novel machine learning approach using Random Forest and Decision Tree classifiers on AOCC data demonstrates the feasibility of predicting optimal topology configurations, establishing a data-driven, explainable framework. While focused on PSO, the methodology is extensible to other swarm-based metaheuristics. Future directions include extending the analysis to higher-dimensional problems, integrating additional XAI techniques, and developing adaptive topology selection systems. Overall, this work shifts optimization research from empirical evaluation toward interpretable, explainable algorithm design, with significant implications for both theoretical understanding and practical applications.
@@
Nitin Gupta: Writing – original draft, Methodology, Experimentation. Bapi Dutta: Review & editing. Anupam Yadav: Writing – review & editing, Supervision.